the heat required to raise the temperature of 30kg of copper from 12 degrees Celsius to 70 degrees Celcius


1
Expert's answer
2022-01-13T13:25:02-0500

"Q=cm(t_2-t_1)=380\cdot 30\cdot(70-12)=670~kJ."